five years ago a man was seven times as old as his son. five years hence, the father will be three times as old as his son. find their present age?
Answer by ankor@dixie-net.com(6693)

let m = man's present age
let s = son's present age
:
Write an equation for each statement:
:
"five years ago a man was seven times as old as his son."
m - 5 = 7(s - 5)
m - 5 = 7s - 35
m = 7s - 35 + 5
m = 7s - 30
:
" five years hence, the father will be three times as old as his son."
m + 5 = 3(s + 5)
m + 5 = 3s + 15
m = 3s + 15 - 5
m = 3s + 10
:
find their present age?
Using the first statement to replace m in the 2nd statement
7s - 30 = 3s + 10
7s - 3s = 10 + 30
4s = 40
s = 10
:
I'll let you find the man's age, check the solutions in both statements.